OpenStudy (anonymous):

Challenge! Ready? What is the solution to the proportion, 3y-8/12= y/5? Show me your work so that I, Lolly, know you are not cheating. THIS IS A PROPORTION; where you cross multiply.

OpenStudy (anonymous):

This is not simple algebra.

OpenStudy (danjs):

is [3y-8]/12 the left side...

OpenStudy (danjs):

\[\frac{ 3y-8 }{ 12 }=\frac{ y }{ 5 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sorry for not saying.

OpenStudy (danjs):

ok, remember to group the numerator if multiple terms...

OpenStudy (danjs):

so if you multiply both sides by a common denominator... in this case (12*5), you,,, nevermind just do the cross multiply thing 5*(3y-8) = 12*y

OpenStudy (danjs):

can you distribute the 5 into that parenthesis term ?

OpenStudy (anonymous):

\[\large \sf 15y-40=12y\]\[\large \sf 3y=40\]\[\large \sf y=\frac{40}{3}\] Right?

OpenStudy (abdullahm):

Hint: \(\sf \Large\frac{a}{b} = \frac{c}{d} \to ad = bc\) And this question as a challenge, and asking for work seems rather suspicious to me...

OpenStudy (anonymous):

MM. You guys are not getting it. : OO

OpenStudy (anonymous):

Are you saying I'm wrong?

OpenStudy (anonymous):

yup

OpenStudy (abdullahm):

Could you please write this: 3y-8/12= y/5 using the equation button or the draw tool?

OpenStudy (anonymous):

I checked my work, and that's what makes the two sides equal...

OpenStudy (anonymous):